# Robust Federated Learning Over the Air: Combating Heavy-Tailed Noise with Median Anchored Clipping

### v1

@@ V Conclusion
In this paper, we proposed a new algorithm, named MAC, which significantly enhances the robustness of OTA FL systems against channel noise that often exhibits heavy-tailed distributions. MAC leverages the median of the global gradient entries as a datum plane, and applies value clipping to truncate the extreme values induced by impulse noise, thereby effectively alleviating the detrimental effects of channel noise while largely preserving the original gradient information. We validated the effectiveness of MAC through convergence analysis and a set of empirical experiments, in which the MAC algorithm demonstrated consistent stability under various noise conditions. The MAC algorithm is effective, low-complex, and straightforward to be implemented, rendering it highly applicable in practical scenarios.

### v4

@@ V Conclusion
In this study, we introduce a novel algorithm, denoted as median anchored clipping(MAC), which markedly enhances the robustness of over-the-air federated learning(OTA FL) systems against channel noise, frequently characterized by heavy-tailed distributions. The MAC algorithm leverages the median of the aggregated global gradient entries as a datum plane and applies value clipping to truncate the extreme values induced by impulse noise, thereby effectively alleviating the detrimental effects of channel noise while largely preserving the original gradient information. The effectiveness of the MAC algorithm was validated through a convergence analysis conducted under non-convex conditions, alongside a series of empirical experiments that demonstrated the algorithm’s consistency across a range of noise conditions. The MAC algorithm is effective, low-complex, and straightforward to implement, rendering it highly applicable in practical scenarios.